OpenStudy (anonymous):

MEDAL will be given! What is the product? 1/5 x 7/8 x 5/6

OpenStudy (anonymous):

1/5x7/8= 7/40

OpenStudy (anonymous):

7/40x5/6= 35/240

OpenStudy (anonymous):

So is the answer 7/40?

OpenStudy (anonymous):

the final answer is 7/48

OpenStudy (anonymous):

in my answer list there is no 7/40

OpenStudy (anonymous):

Yes I know it is 35/240 but that is not in the answer list

OpenStudy (anonymous):

I want an answer.

OpenStudy (anonymous):

ummmm

OpenStudy (anonymous):

I got 7/48 what are the answer choices?

OpenStudy (anonymous):

thanks for the answer, ill give a medal if it is right

OpenStudy (anonymous):

um

OpenStudy (anonymous):

can i ask more?

OpenStudy (anonymous):

that was correct.

OpenStudy (anonymous):

can i ask more now?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

If its algebra 1 I can help